Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 14     Entry time: 07 Feb 2025, 05:56
Author: Peter Zalden 
Group: FXE 
Subject: Shift summary 

Preparation for user expt #7944

X-ray delivery

  • since pulses (10 Hz) at 9.3 keV, 3 mJ
  • Pointing stability 55 um RMS in the horizontal plane on FXE PI, which is lower as compared to earlier values

Achievements/Observations

  • Aligned xrays into FXE vacuum/cryo setup
  • Performed ice depositions in-situ and characterized by Xray

Handover

  • Beam transport to SPB and 11 keV recovered
  • We left Hirex inserted during FXE operation, but closed the GH2 shutter. It is still closed.

Issues

  • XTD2 crl lens3x motor moving in wrong direction and already past all limits. Created ticket with EEE (no DOC available) #216784. Lens3x is not needed for the ongoing experiment, but the situation is potentially dangerous.
ELOG V3.1.4-7c3fd00